In an A.P, the 'p'th term is q and the (p+q)th term is 0, then common difference is:
A. -p
B. q
C. -1
D. 0
 

Grade:12th pass

1 Answers

Vikas TU
4 years ago
Dear student 
Pth term = a+ (p-1)d = q 
a+ (pd-q-d) =0...........1
(p+q )th term = 0 = a+ (p+q-1)d …..........2
Solving 1 and 2 
-2qd=q
d = -1/2